OpenStudy (anonymous):

WHY ISN't anyone helping me??? Please help me I have a final tomorrow :( (3/2x)-(5/X^2-5x)+(x/x-5)

OpenStudy (anonymous):

\[\frac{3}{2x}-\frac{5}{x^2-5x}+\frac{x}{x-5}\] like that?

OpenStudy (anonymous):

Yes thank you.

OpenStudy (anonymous):

you have to set all the denominators the same

OpenStudy (anonymous):

you need a common denominator since \(x^2-5x=x(x-5)\) you can use as a denominator \[2x(x-5)\]

OpenStudy (anonymous):

do I multiply the numerator and denom by 2x(x-5)?

OpenStudy (anonymous):

the just like with numbers \[\frac{3(x-5)-5(2)+x(2x)}{2x(x-5)}\]

OpenStudy (anonymous):

you only have to multiply by what is missing from the denominator, just like with numbers \[\frac{3}{2x}=\frac{3(x-5)}{2x(x-5)}\] \[-\frac{5}{x(x-5)}=-\frac{5\times 2}{2x(x-5)}\] \[\frac{x}{x-5}=\frac{x\times 2x}{2x(x-5)}\]

OpenStudy (anonymous):

like if you wanted to use 24 for the denominator to add \(\frac{5}{6}+\frac{7}{8}\) you would need \[\frac{5}{6}=\frac{5\times 4}{24}\] \[\frac{7}{8}=\frac{7\times 3}{24}\] it is the same idea exactly

OpenStudy (anonymous):

ok satelite I gt 3(x-5/2x(x-5)-5x2/2x(x-5)+X x 2x/2x(x-5)

OpenStudy (anonymous):

when I distribute I get 3x-15/2x^2-10x

OpenStudy (anonymous):

put them all over the same denominator, then multiply out in the numerator and combine like terms you have a mistake somewhere because there is a \(2x^2\) in the numerator

OpenStudy (anonymous):

oh Okay but how do I multiply x(x-5) times 2

OpenStudy (anonymous):

\[3(x-5)-5\times 2+2x^2=3x-15-10+x^2\] \[=x^2+3x-25\]

OpenStudy (anonymous):

Oh I see..

OpenStudy (anonymous):

\(x(x-5)\times 2=2x(x-5)=2x^2-10x\) to answer your second question

OpenStudy (anonymous):

oh ok gotcha!

OpenStudy (anonymous):

Thanks a lot!